# Sven Lilge and Timothy D. Barfoot Incorporating Control Inputs in Continuous-Time Gaussian Process State Estimation for Robotics

### v2

@@ 4.1 Experimental Setup
For the experiments, we are using the tendon-driven continuum robot (TDCR) prototype and sensor setup presented in [5] (see Fig. 5, left). This prototype features two controllable segments, with tendons routed parallel to the backbone through spacer disks and attached at the end of each segment. The bending of both segments can be individually actuated by pulling and releasing the tendons. For sensing, we consider a single six-degree-of-freedom pose measurement at the robot’s tip using an electromagnetic tracking sensor (Aurora v3, Northern Digital Inc., Canada). Ground-truth is obtained using a coordinate measurement arm with an attached laser probe (FARO Edge with FARO Laser Line Probe HD, FARO Technologies Inc., USA). This measurement arm allows us to capture a detailed laser scan of the robot’s shape in the form of a point cloud, from which the discrete pose of each spacer disk can be reconstructed. Six simple bending configurations of the continuum robot were exclusively recorded to calibrate the sensor and ground-truth frames. Using the same procedure outlined in [5], the remaining calibration error at the tip of the robot results in mm.
@@ 4.2 Evaluation
For state estimation, we now wish to compare the conventional prior formulation from [5] with the newly proposed one considering inputs. As discussed above, there are two possibilities to incorporate known inputs. One can either consider known input strains or known distributed actuation loads . Either of these possible outputs can, for instance, be provided by another open-loop continuum robot model.
@@
Both the traditional prior from [5] as well as the newly proposed prior considering inputs are fused with a single pose measurement at the tip of the robot. As an additional degenerate case, we further consider the scenario in which only the robot tip’s position can be measured. For every case, we compute the RMSE and maximum errors for the robot’s position and orientation along its shape, considering the discrete ground-truth poses at each disk.
@@
4.3 Results
The results for each of the nine configurations, as well as their average, are reported in Tab. 2. In each case, the state estimation problem could be solved in 5-10 ms. Fig. 5, middle and right, additionally showcases the state estimation results for the first configuration, both when considering a full pose measurement and a position-only measurement of the continuum robot’s tip. It can be seen that due to the ability to retain the local shape of the prior, our proposed method outperforms the conventional GP formulation in every case, with two exceptions. In these cases, relatively high external forces were applied to the continuum robot, which resulted in shapes that significantly differ from the utilized prior, making it less helpful. Nevertheless, the results in these cases remain largely comparable to those obtained by the baseline method.
In the remaining configurations, the accuracy of the proposed method is significantly higher than that of the baseline method, especially in cases where only the position of the robot’s tip is measured. In these instances, the baseline method’s shape estimates largely deviate from the ground-truth, with worst-case maximum errors exceeding 3 cm and 45∘. This further underscores the proposed method’s effectiveness in situations with sparse sensing, as the additional inputs contribute to a more informed prior, resulting in more a accurate state estimation. This is also apparent in the example shown in Fig. 5.
@@ 5 Conclusion
In this work, we propose an extension to the WNOA GP prior used in continuous-time batch state estimation. Our new formulation allows us to incorporate both known velocity and acceleration inputs, as long as they can be expressed in the robot’s body frame independently of the robot state. We specifically present the case of piecewise-linear inputs, for which closed-form approximations can be obtained to represent motions in . Experiments showcase the use cases of the method for both mobile robot trajectory and continuum robot shape estimation. It is shown that considering inputs in the GP prior is particularly helpful to maintain the local shape of the trajectory between discrete state observations, specifically when the motion differs from the constant-velocity assumption of the traditional WNOA GP prior. Future work will focus on further evaluating the approach on mobile robots with more aggressive motions, such as present in drone racing, in which cases the traditional WNOA GP prior is less accurate. Further, it will be investigated whether inputs can also be considered in the WNOJ GP motion prior as an alternative.

### v3

@@ 4.1 Experimental Setup
For the experiments, we are using the tendon-driven continuum robot (TDCR) prototype and sensor setup presented in [22] (see Fig. 6). This prototype features two controllable segments, with tendons routed parallel to the backbone through spacer disks and attached at the end of each segment. The bending of both segments can be individually actuated by pulling and releasing the tendons. For sensing, we consider a single six-degree-of-freedom pose measurement at the robot’s tip using an electromagnetic tracking sensor (Aurora v3, Northern Digital Inc., Canada). Ground-truth is obtained using a coordinate measurement arm with an attached laser probe (FARO Edge with FARO Laser Line Probe HD, FARO Technologies Inc., USA). This measurement arm allows us to capture a detailed laser scan of the robot’s shape in the form of a point cloud, from which the discrete pose of each spacer disk can be reconstructed. Six simple bending configurations of the continuum robot were exclusively recorded to calibrate the sensor and ground-truth frames. Using the same procedure outlined in [22], the remaining calibration error at the tip of the robot results in mm.
@@ 4.2 Evaluation
For state estimation, we now wish to compare the conventional prior formulation from [22] with the newly proposed one considering inputs. As discussed above, there are two possibilities to incorporate known inputs. One can either consider known input strains or known distributed actuation loads . Either of these possible inputs can, for instance, be provided by another open-loop continuum robot model.
@@
Both the traditional prior from [22] as well as the newly proposed prior considering inputs are fused with a single pose measurement at the tip of the robot. As an additional degenerate case, we further consider the scenario in which only the robot tip’s position can be measured. For every case, we compute the RMSE and maximum errors for the robot’s position and orientation along its shape, considering the discrete ground-truth poses at each disk. Hyperparameters are tuned empirically on representative continuum robot configurations. For both methods, the prior covariance is set to , while the pose measurement covariance is .
4.3 Results
The results for each of the nine configurations, as well as their average, are reported in Tab. 3. In each case, the state estimation problem could be solved in 5-10 ms. Fig. 7 additionally showcases the state estimation results for the first configuration, both when considering a full pose measurement and a position-only measurement of the continuum robot’s tip. It can be seen that due to the ability to retain the local shape of the prior, our proposed method outperforms the conventional GP formulation in every case, with two exceptions. In these cases, relatively high external forces were applied to the continuum robot, which resulted in shapes that significantly differ from the utilized prior, making it less helpful. Nevertheless, the results in these cases remain largely comparable to those obtained by the baseline method.
@@ 4.3 Results
In the remaining configurations, the accuracy of the proposed method is significantly higher than that of the baseline method, especially in cases where only the position of the robot’s tip is measured. In these instances, the baseline method’s shape estimates largely deviate from the ground-truth, with worst-case maximum errors exceeding 3 cm and 55∘. This further underscores the proposed method’s effectiveness in situations with sparse sensing, as the additional inputs contribute to a more informed prior, resulting in more a accurate state estimation. This is also apparent in the example shown in Fig. 7.
@@ 5 Conclusion
In this work, we propose an extension to the WNOA GP prior used in continuous-time batch state estimation. Our new formulation allows us to incorporate both known velocity and acceleration inputs, as long as they can be expressed in the robot’s body frame independently of the robot state. We specifically present the case of piecewise-linear inputs, for which closed-form approximations can be obtained to represent motions in . Experiments showcase the use cases of the method for both mobile robot trajectory and continuum robot shape estimation. It is shown that considering inputs in the GP prior is particularly helpful to maintain the local shape of the trajectory between discrete state observations, specifically when the motion differs from the constant-velocity assumption of the traditional WNOA GP prior.
The main advantage of the proposed approach is the introduction of more informed priors, potentially reducing the number of measurements and estimation nodes required to achieve accurate state estimates. This reduction can significantly decrease the computation time for state estimation problems, as shown during mobile robot experiments. Furthermore, as shown during the continuum robot experiments, the approach can enable more accurate estimations in scenarios with limited sensing, as the inclusion of control inputs adds valuable information to the prior, compensating for the lack of measurements.
Nevertheless, the proposed method has some limitations. Incorporating control inputs increases the complexity of computing the prior and posterior mean and covariances. As a result, both the MAP optimization process and state querying require more computation time compared to the baseline method. Favourable computation times can only be achieved when using less frequent estimation nodes, where the proposed approach can still maintain estimation accuracy. Another limitation arises when the control inputs exhibit discontinuities, which persist in the posterior distribution and reduce its smoothness. While this characteristic may be advantageous in certain scenarios, it can negatively impact accuracy when noisy measurements are directly used as inputs. Furthermore, for mobile robots, the presented experiments did not demonstrate a significant improvement in accuracy when incorporating velocities as inputs rather than as measurements, particularly with frequent discrete estimation nodes. This may be due to the fact that the traditional WNOA prior already models the robot’s motion quite effectively, assuming constant velocity between discrete time steps.
Future work will focus on evaluating the proposed approach in scenarios involving more aggressive motions, such as drone racing, where the traditional WNOA GP prior is less effective. This will help identify situations in which incorporating control inputs into GP priors is particularly advantageous. Additionally, we aim to investigate whether similar benefits can be achieved by incorporating inputs into the WNOJ GP motion prior as an alternative.